Write an equation of the line containing the specified point and parallel to the indicated line.
(−7,0),5x+2y=6

Answer by nerdybill(7384)

(−7,0),5x+2y=6
.
First, determine slope of:
5x+2y=6
.
Do this by putting it into "slope-intercept" form:
5x+2y=6
2y = -5x+6
y = (-5/2)x + 3
slope is -5/2
.
since our line is PARALLEL it must have ths same slope.
taking our slope (-5/2) and the point (-7,0) plug it into the "point-slope" form:
y - y1 = m(x - x1)
y - 0 = (-5/2)(x - (-7))
y = (-5/2)(x + 7)
y = (-5/2)x + (-5/2)7
y = (-5/2)x + (-35/2)
y = (-5/2)x - 35/2 (this is what they're looking for)